# Volumetric Functional Maps

### v2

@@ 5.1. Volumetric Functional Maps
Figure 5. Left: average geodesic error curves on the pairs from Su et al. (Su et al., 2019) resulting from the application of FMaps (Nogneng and Ovsjanikov, 2017), ZoomOut (Melzi et al., 2019b), and Orthoprods (Maggioli et al., 2021). Right: The execution time for the three methods plotted against the number of vertices in the source mesh.
Figure 6. Coordinate transfer between two volumetric shapes using the Orthoprods basis. The normalized coordinates are treated as RGB channels for visualization (first column). For further validation, the transferred coordinates are also used for generating an error sensitive procedural texture, which is visualized on different slices of the volume (second to fourth columns).
Figure 7. Number of flipped tetrahedra obtained by transferring the mesh connectivity considering a growing number of basis functions. From left to right: coordinate transfer using the LBO basis; coordinate transfer using the CMH basis; coordinate extrapolation using the LBO basis; coordinate extrapolation using the CMH basis. The smaller colored lines show the trend for each pair of shapes in our reference dataset (Su et al., 2019). The bold black line is the average trend across the entire dataset. The two pairs that take longer to converge with the LBO basis functions (yellow and orange lines) are the ones where source and target differ the most and are strongly non-isometric. This explains why using only the intrinsic information from the LBO eigenbasis makes the method converge slowly and adding extrinsic information from CMH aligns the pairs with the other results.
@@
To validate our prototype and assess its ability to devise a high-quality volume map, we compare the correspondences we compute against the ground truth. Unlike the surface setting, where in recent years there has been a constantly growing amount of public data useful for validation, almost no volumetric datasets are available for our specific task. We consider a set of 40 pairs of tetrahedral meshes with identical connectivity, which thus define an injective piece-wise linear map as described in the beginning of Section 2.1. This dataset, that was originally proposed by (Su et al., 2019) and then used to validate (Du et al., 2020), covers diverse classes of objects, such as: animals, humanoids, hands, and a variety of man-made objects like glasses, lamps, laptops and scissors. For each pair of objects we compute a volumetric functional map and extract a point-to-point correspondence as described in Section 4.1. Then, for each vertex $v$ we measure the geodesic error, which is the geodesic distance $d_{\mathcal{N}}(\pi_{\mathrm{gt}}(v),\pi(v))$ between its mapped point $\pi(v)$ and its ground truth counterpart $\pi_{\mathrm{gt}}(v)$. In Figure 5 we show the cumulative curves of the geodesic error obtained with different functional map estimation algorithms, as well as their runtime in relation to the vertex count. We also provide the average geodesic error (AGE) in the legend of the figure. For the comparison, we use the well-established algorithm based on descriptors preservation (FMap) (Nogneng and Ovsjanikov, 2017) and ZoomOut (Melzi et al., 2019b), using both the standard and the fast implementation. For the fast implementation, we use the surface points as samples to validate the claim in Section 4.2 that the surface encodes information about the enclosed volume. We also test our framework replacing the LBO eigenbasis with the orthogonalized eigenproducts basis (Orthoprods) proposed by Maggioli et al. (2021), but instead of using the approach described by the authors for the functional map estimation, we plug the basis into the ZoomOut pipeline. As expected, the extended Orthoprods basis has a higher descriptive power, but this comes with an additional computational cost for aligning the larger bases. In contrast, ZoomOut achieves the best trade-off between accuracy and runtime. Overall, our results demonstrate that the volumetric functional map framework produces high quality maps, matching the performance obtained in the surface setting. This can be further appreciated in the signal transfer example from Figure 6, where we transfer the coordinates function between two meshes with the Orthoprods bases and use the resulting values to compute an error sensitive procedural texture (Maggioli et al., 2022).
@@
Table 1. Average percentage of flips for the methods from Figure 7 at varying of the basis size. The last column also reports the average time (in seconds) required for the computation of the largest basis considered.
@@
Figure 8. Left: two of the 1405 failure cases of the untangling algorithm proposed by Garanzha et al. (2021), which emerged from the challenging large scale validation described in (Nigolian et al., 2023, §6.3). Right: computing an initial connectivity transfer with our tool to warm start the same algorithm allows to retain a fully bijective map, thus improving the overall robustness.
Figure 9. From left to right: a tetrahedral mesh; the external surface of the target domain; the ground truth connectivity transfer, obtained with the volumetric ARAP implementation in Cinolib (Livesu, 2019); the connectivity transfer obtained by extrapolating the interior vertices from the surface coordinates; the connectivity transfer obtained by transferring the coordinates with the functional map. Each tetrahedron $t$ is colored according to the geometric distortion of the map, measured as $|1-\det(J_{t})|$, where $J_{t}$ is the Jacobian of the affine transformation mapping $t$ from the source to the target shape. Both the surface extrapolation and volumetric transfer approaches yield maps that are visually identical with the ground truth. All maps shown in this figure – both the ground truth and the ones that we computed – are fully injective (i.e., $\forall t,\det(J_{t}$)¿0).
5.2. Connectivity Transfer
@@
Representing volume maps using two tetrahedral meshes that share the same connectivity but have different vertex coordinates is a common choice for many practical applications involving digital fabrication (Liu et al., 2024; Zhang et al., 2022; Etienne et al., 2019), medicine (Abulnaga et al., 2021) and hexahedral meshing (Pietroni et al., 2022; Brückler and Campen, 2023). As detailed in Section 4.2, our framework can be used to transfer the connectivity of a given tetrahedral mesh onto a target domain, either by using a volumetric functional map (§4.2 Functional Maps for Connectivity Transfer) or by extrapolating the connectivity from its outer surface (§4.2 Spectral Representation of Volumes).
@@
If the whole spectrum is considered, the LBO basis spans the entire functional space of the mesh and the functional map yields one-to-one correspondences (Rodola et al., 2017). Computing the full spectrum is computationally expensive, hence only a fraction of the eigenfunctions are typically computed. We test our prototypes on the dataset of 40 meshes released by Su  et al. (2019) and empirically verify that the amount of inverted elements introduced in the piece-wise map gracefully decays to zero for growing numbers of eigenfunctions (Figure 7). On average, already using $5\%$ of the spectrum the number of flipped tetrahedra is below $5\%$. It goes below $1\%$ if $20\%$ of the spectrum is considered (Table 1). We consider both the LBO spectrum and its extension equipped with three additional functions that encode the extrinsic embedding (CMH). As shown in (Marin et al., 2019; Melzi et al., 2020) CMH is better suited to transfer discrete signals such as the surface mesh connectivity. This advantage also manifests on volumes, as reflected in the second and fourth columns in Figure 7 and in the second and fourth lines of Table 1. For efficiency, the average running times for computing the larger bases are reported in the last column of the same table. As can be noticed the extrapolation algorithm is roughly 2×2$\times$2 × faster than volumetric transfer, because in this case only the volumetric spectrum of a single mesh must be computed. Interestingly, if the mesh resolution is sufficiently low and the computation of a higher percentage of the spectrum is computationally feasible, we can directly devise a fully injective piecewise linear map, as shown in Figure 9, where $25\%$ of the spectrum is considered.
When computational resources are limited and the full spectrum cannot be computed, volumetric functional mapping still provides a valuable initial solution for untangling algorithms such as (Garanzha et al., 2021; Du et al., 2020), which are then required to remove the few inverted tetrahedra present in the map. This is a standard procedure for the generation of injective linear maps, where mesh untangling is typically bootstrapped using the Tutte embedding, which is known to flip volumetric cells (Alexa, 2023). Volumetric functional map provides a better initial solution than Tutte. To prove this, we considered the 1405 failures produced in the large-scale validation described in (Nigolian et al., 2023, §6.3). Substituting the initial solution provided by Tutte with the one produced with our framework, we were able to produce fully bijective maps in the 85% of their failure cases using both LBO and CMH basis functions, thus dramatically increasing the success rate of this pipeline. Two representative examples are shown in Figure 8.
5.3. Segmentation Transfer
@@
The functional maps framework proved to be particularly successful in the area of statistical shape analysis (Stegmann and Gomez, 2002). In particular, the study of medical data primarily benefits from accurate solutions for analyzing large collections of shapes (Heimann and Meinzer, 2009), and the literature proves that the functional maps framework can be successfully applied for medical applications (Melzi et al., 2016; Magnet et al., 2023; Maccarone et al., 2024). A standard approach for analyzing collections of shapes is to determine or build a template surface which acts as a median shape for comparing the entire dataset (Huang et al., 2019). However, existing approach are limited to the analysis of surfaces, while often medical data are better represented as 3D volumes, such in the case of MRI scans (Rinck, 2019).
@@
In this setting, our approach allows to extend existing methodologies to deal with volumetric representations. To prove this, we select two brain shapes from the MedShapeNet dataset (Li et al., 2024) and a population-averaged template (Evans et al., 1993). Using TetGen (Hang, 2015), we compute the tetrahedral meshes representing the volumes of the brains, and segment the template using a reference segmentation (Fischl et al., 2002). We then compute a volumetric functional map from the template to each shape, extract the point-to-point correspondence and transfer the segmentation. The results are reported in Figure 10, where we show the segmentation on both the surface and on a slice of the volume. We can see that, despite the noisy segmentation in the central section of both the source and the target domains, our tool can faithfully transfer the segmentation preserving its overall structure.
@@
Figure 10. A segmentation of a template brain transferred to two other brains using a correspondence computed with our volumetric functional maps framework. The shapes and the segmentation are from MedShapeNet (Li et al., 2024).
5.4. Shape Matching
@@
Besides the ability to transfer volumetric signals, we also study to what extent incorporating volumetric information onto the shape matching problem improves accuracy over existing surface-based functional maps pipelines.
@@
We consider both the outer surfaces of our reference ground truth volumetric dataset (which we will refer to as VOL) (Su et al., 2019) and well established datasets for shape matching, such as SHREC’19 (Melzi et al., 2019a), SHREC’20 (Dyke et al., 2020), and TOPKIDS (Lähner et al., 2016), which we all tetrahedralize using fTetWild (Hu et al., 2020). Overall, these datasets stress shape matching algorithms in different ways, exhibiting geometric and topological noise (SHREC’19), strongly non-isometric matches (SHREC’20) and topological changes (SHREC’20 and TOPKIDS). For SHREC’20, we discard two partial meshes because the inside/outside system of fTetWild was not able to correctly reproduce a volumetric mesh that allows a meaningful comparison with surface methods. More discussion about the tetrahedralization of the surfaces is reported in Appendix E.
Figure 11. Cumulative plot of the geodesic error obtained with the standard functional maps framework (blue) on surfaces and our volumetric functional maps approach (orange) on four different datasets. The results are averaged across each dataset.
Dataset Basis AGE Surf. AGE Vol. Succ. Rate Slowdown
@@
Table 2. The specific functional maps implementation used for each dataset. We also report the geodesic error averaged across the entire dataset (AGE) of both the surface and volumetric approach, as well as the percentage of single pairs where our approach obtains a better AGE (Succ. Rate) and the average slowdown factor induced by using our method. As can be noticed, computational overhead heavily correlates with mesh growth, which we measure as the average ratio between the number of vertices in the volume and the surface meshes and we report between parentheses.
@@
For validating the approach discussed in Section 4.3, we compare it with the standard functional maps framework for surfaces on each dataset presented in the previous paragraph. In order to prove that our volumetric framework can be seamlessly integrated with existing extensions of functional maps, we also test it with different sets of bases, such as the Coordinates Manifold Harmonics (CMH) (Marin et al., 2019) and the orthogonalized eigenproducts (Orthoprods) (Maggioli et al., 2021). In all cases, we align the bases using the method based on descriptors preservation (Nogneng and Ovsjanikov, 2017) and ZoomOut (Melzi et al., 2019b). For a fair comparison, we use the same basis for both the surface and volumetric approach. Details can be found in Table 2
Figure 12. The tool we use to tetrahedralize surface matching datasets (Hu et al., 2020) is robust, but occasionally introduces topological noise, closing unwanted handles that alter the mapping task (closeup). The LBO eigenbasis is not robust against such defects. Conversely, the CMH basis is robuster and greatly improves the mapping quality. Both the camel and the cow are from the SHREC’20 dataset.
@@
The table also reports the dataset averaged geodesic error (AGE) for both the frameworks and the percetange of single pairs for which our approach obtains a better AGE (Success Rate), as well as the slowdown factor induced by the introduction of the volume and the ratio between the vertex count of the tetrahedral and the surface meshes (between parentheses). More detailed results about the comparison can be found in Figure 11, where we show the cumulative geodesic error curves averaged across each dataset and their corresponding area under the curve (AUC). Interestingly, our approach is consistently much more accurate than its surface counterpart, suggesting that incorporating volumetric information into the matching process is always a rewarding choice when a volumetric discretization is available. The increase in accuracy comes at the cost of some additional computational effort, primarily due to the introduction of the internal vertices. By inspecting the last column of Table 2, we can see that there is a close relationship between the slowdown and the number of vertices introduced, suggesting that the volumetric pipeline is not significantly slower per se, but suffers from the overhead induced by the additional vertices needed to discretize the volume.
SHREC’20 is the only dataset where the surface-only method occasionally obtains better results (in the 7.27% of the cases). Based on our understanding, these sporadic cases in which surface-based methods surpass our volume pipeline depend on the basis alignment algorithm we use (Nogneng and Ovsjanikov, 2017), which is known to not be robust under strong non-isometric deformations and topological changes, which occasionally occurred when tetrahedralizing our surface-based reference datasets. This behavior can be partially compensated by using more informative bases such as CMH, which provide a better tool for representing extrinsic information. This can be better appreciated in the example from Figure 12, where we show a surface coordinate transfer in a pair from the SHREC’20 dataset with our volumetric approach using the standard LBO eigenbasis and the CMH basis. However, integrating more robust basis alignment methods is certainly worth of future investigation.
6. Conclusions
@@
We have presented the first volumetric extension of the functional maps pipeline, showcasing practical applications that build on top of this connection to transfer both discrete and continuous signals for solid texturing, connectivity transfer and segmentation transfer for medical applications, as well as to improve the accuracy of shape matching tasks on surfaces. We believe that our practical contributions just scratch the surface of this problem and foresee interesting avenues for future research.
@@
Our results on connectivity transfer show great potential but are far from being perfect. Indeed, achieving a fully injective map is technically possible, but considering a large enough portion of the spectrum may easily become computationally infeasible for applications. Devising smarter techniques to retain more precision for fixed spectrum size is an interesting direction worth investigating.
The higher precision in shape matching tasks comes at the cost of an additional computational effort, because the most expensive steps in the pipeline (i.e., the eigendecomposition and the alignment of the bases) suffer from the increased vertex count of the volume mesh. This limitation, which is intrinsic of our approach, could be mitigated by integrating scalable algorithms for computing and aligning the bases specifically designed for volume meshes.
@@
First and foremost, a theoretical study of what properties are invariant under a volume functional map must be carried out. Despite our empirical studies show that volumetric maps are more accurate than their surface counterparts, a more profound theoretical analysis of this behavior would provide useful insights for the development of new solutions.
Secondly, the success of the functional maps framework was also related to its high flexibility and easy extensibility. A variety of solutions have been devised to extend the LBO eigenbasis, refine the map, and integrate the framework in machine learning pipelines. In this work, we just scratched the surface of this line of research, showing that some of the existing extensions can be seamlessly integrated into the volumetric framework. However, other works would require deeper analyses to generalize to tetrahedral meshes.
Finally, we emphasize the importance of releasing new volumetric datasets to support ongoing research on volume mapping in general, and on spectral volume methods in particular. As discussed in Section 5.4 surface-based datasets that are well established in the functional map community hardly extend to volumes because of the difficulty to produce tetrahedral meshes of acceptable quality, due to the presence of both topological and geometric noise, (e.g., large missing parts, overlaps, multiple connected components, non-manifold configurations and self-intersections). While robust volume meshing methods such as (Hu et al., 2020; Diazzi and Attene, 2021) can handle intersections with ease, determining semantically plausible separations between the inside and outside portions of the volume in the presence of large gaps remains an open meshing challenge (Figure 15). Recently, a large scale dataset to support ongoing research on piece-wise volume mappings to convex or star-shaped domains has been released (Cherchi and Livesu, 2023). Having similar contributions to support spectral methods would greatly help practitioners and ultimately foster more research in this emerging field.

### v3

@@ 5.1 Volumetric Functional Maps
Figure 4: Left: average geodesic error curves on the pairs from Su et al. [113] resulting from the application of FMaps [86], ZoomOut [79], and Orthoprods [70]. Right: runtime for the three methods plotted against the number of vertices in the source mesh.
@@
We validate our prototype and assess its ability to devise high-quality volume maps by comparing our correspondences against the ground truth. Unlike the surface setting, where a constantly growing amount of public data has become available for validation, almost no volumetric datasets are available for this specific task. We consider a set of 40 pairs of volume meshes with identical connectivity, which thus define an injective piece-wise linear map as described in Sec. 2.1. This dataset, originally proposed by [113] and then used to validate [28], covers diverse classes of objects, such as: animals, humanoids, hands, and a variety of man-made objects like glasses, lamps, laptops and scissors. For each pair of objects, we compute a volumetric functional map and extract a point-to-point correspondence as described in Sec. 4. Then, for each vertex $v$, we measure the geodesic error as the geodesic distance $d_{\mathcal{N}}(\pi_{\mathrm{gt}}(v),\pi(v))$ between its mapped point $\pi(v)$ and its ground truth counterpart $\pi_{\mathrm{gt}}(v)$. In Fig. 4 we show the cumulative curves of the geodesic error obtained with different functional map estimation algorithms, as well as their runtime in relation to the vertex count. We also provide the average geodesic error (AGE) in the legend of the figure. For comparisons, we use the well-established algorithm based on descriptor preservation (FMap) [86] and ZoomOut [79], using both the standard and the fast implementation. For the fast implementation, we use the surface points as samples to validate the claim in Sec. 4.1 that the surface encodes information about the enclosed volume. We also test our framework replacing the LBO eigenbasis with the orthogonalized eigenproducts basis (Orthoprods) proposed by Maggioli et al. [70], but instead of using the approach described by the authors for the functional map estimation, we plug the basis into the ZoomOut pipeline. As expected, the extended Orthoprods basis has higher descriptive power, but this comes with additional computational cost for aligning the larger bases. In contrast, ZoomOut achieves the best trade-off between accuracy and runtime. Overall, our results demonstrate that the volumetric functional map framework produces high-quality maps, matching the performance obtained in the surface setting. This can be further appreciated in the signal transfer example from Fig. 5, where we transfer the coordinates function between two meshes with the Orthoprods bases and use the resulting values to compute an error-sensitive procedural texture [68]. All values reported in this subsection correspond to volumetric extensions of the considered metrics. A more detailed evaluation of the maps is provided in Appendix E.
Figure 5: Coordinate transfer between two volumetric shapes using the Orthoprods basis. The normalized coordinates are treated as RGB channels for visualization (first column). For further validation, the transferred coordinates are also used for generating an error-sensitive procedural texture, which is visualized on different slices of the volume (second to fourth columns).
@@
Table 1: Average percentage of flips using the methods described in Sec. 4.1 at varying the basis size. The average runtime (in seconds) to compute the largest basis considered is also reported.
@@
Figure 6: Left: two of the 1405 failure cases of the untangling algorithm proposed by Garanzha et al. [36], which emerged from the challenging large scale validation described in [83, §6.3]. Right: computing an initial connectivity transfer with our tool to warm start the same algorithm allows to retain a fully bijective map, thus improving the overall robustness.
5.2 Connectivity Transfer
@@
Representing volume maps using two tetrahedral meshes that share the same connectivity but have different vertex coordinates is a common choice for many practical applications involving digital fabrication [59, 122, 30], medicine [2] and hexahedral meshing [91, 12]. As detailed in Sec. 4.1, our framework can be used to transfer the connectivity of a given volume mesh onto a target domain, either by using a functional map to transfer the vertex coordinates (§4.1 Functional Connectivity Transfer) or by extrapolating the connectivity from its outer surface (§4.1 Spectral Extrapolation of Coordinates).
@@
If the whole spectrum is considered, the LBO basis spans the entire functional space of the mesh and the functional map is one-to-one [102]. Computing the full spectrum is computationally expensive, hence only a fraction of the eigenfunctions are typically computed. We test our prototypes on the dataset of 40 meshes released by Su et al. [113] and empirically verify that the amount of inverted elements introduced in the map gracefully decays to zero for growing numbers of eigenfunctions (see Appendix F for a detailed analysis). On average, already using $5\%$ of the spectrum the number of flipped cells falls below $5\%$. It consistently decreases as basis grows, and goes below $1\%$ if $20\%$ of the spectrum is considered (Tab. 1), reaching full injectvity when 25% of the spectrum is used (Fig. 7). We consider both the LBO spectrum and its extension equipped with three additional functions that encode the extrinsic embedding (CMH). As shown in [73, 76] CMH is better suited to transfer discrete signals such as the surface mesh connectivity. This advantage also manifests on volumes, as reflected in the second and fourth rows of Tab. 1. For efficiency, the average running times for computing the larger bases are reported in the last column of the same table. As can be noticed the extrapolation algorithm is roughly $2\times$ faster than volumetric transfer, because in this case only the volumetric spectrum of a single mesh must be computed. Note that even using a small fraction of the spectrum, volumetric functional mapping still provides a valuable initial solution for untangling algorithms such as [36, 28], which are then required to remove the remaining inverted tetrahedra from the map. This is a standard procedure for generating injective linear maps, where mesh untangling is typically bootstrapped using the Tutte 3D embedding [41], which is known to fail on volumes [8]. Our tool provides a better initial guess than Tutte. To prove this, we considered the 1405 failures produced in the large-scale validation described in [83, §6.3]. Substituting the initial guess provided by Tutte with the one produced with our tool, we were able to obtain one-to-one maps in the 85% of their failure cases, using both LBO and CMH basis functions, dramatically increasing the success rate of this pipeline (Fig. 6).
Figure 7: From left to right: tetrahedral mesh; surface of the target domain; ground truth connectivity transfer; connectivity transfer obtained by extrapolating the interior vertices from the surface coordinates; connectivity transfer obtained by transferring the coordinates with the functional map. Each tetrahedron $t$ is colored according to the geometric distortion of the map, measured as $|1-\det(J_{t})|$, where $J_{t}$ is the Jacobian of the affine transformation mapping $t$ from the source to the target. All maps shown in this figure are fully injective (i.e., $\forall t,\det(J_{t})>0$).
5.3 Segmentation Transfer
@@
Statistical shape analysis for medical data primarily benefits from accurate solutions to analyze large collections of shapes [40, 111]. The literature shows that the functional maps framework is a prominent tool for medical applications [78, 71, 67]. A standard approach to analyze collections of shapes is to determine or build a template surface which acts as a median shape to compare the entire dataset [44]. However, existing approach are limited to the analysis of surfaces, while often medical data are better represented as 3D volumes (e.g., MRI scans [101]).
@@
Our approach allows to extend existing methodologies to deal with volumetric data. To prove this, we select two brain shapes from the MedShapeNet dataset [55] and a population-averaged template [31]. Using TetGen [39], we compute the tetrahedral meshes representing the volumes of the brains, and segment the template using a reference segmentation [34]. We then compute a volumetric functional map from the template to each shape, extract the point-to-point correspondence and transfer the segmentation. From the results in Fig. 8, we can see that, despite the noisy segmentation in the central section, our tool can faithfully transfer the segmentation preserving its overall structure.
@@
Figure 8: A segmentation of a template brain transferred to two other brains using a correspondence computed with our volumetric functional maps framework. The shapes and the segmentation are from MedShapeNet [55].
5.4 Shape Matching
@@
Here we study to what extent incorporating volumetric information into the shape matching problem improves accuracy over existing surface-based functional map pipelines.
@@
We consider both the outer surfaces of our reference ground truth volumetric dataset (denoted as VOL) [113] and well established shape matching datasets, such as SHREC’19 [77], SHREC’20 [29], and TOPKIDS [50], which we all tetrahedralize using fTetWild [43]. These datasets stress shape matching algorithms in different ways, exhibiting geometric and topological noise (SHREC’19), strongly non-isometric matches (SHREC’20) and topological changes (SHREC’20 and TOPKIDS). For SHREC’20, we discard two meshes because the inside/outside system of fTetWild does not reproduce a meaningful volumetric mesh for the comparison. We refer to Appendix G for more details on the data.
LBO CMH Orthoprods
Dataset AGE Surf. AGE Vol. Succ. Rate AGE Surf. AGE Vol. Succ. Rate AGE Surf. AGE Vol. Succ. Rate Slowdown Vertex Ratio
@@
Table 2: Results of each functional maps implementation on each dataset. We report geodesic error averaged across the dataset (AGE) and percentage of single pairs where our approach obtains a better AGE (Succ. Rate). Computational overhead strongly correlates with mesh growth (right). We report: average slowdown induced by volumes and average vertex ratio between surfaces and volumes.
@@
We validate the approach in Sec. 4.2 by comparing it with the standard surface-based pipeline on each dataset. To prove that our volumetric framework can be seamlessly integrated with existing extensions of functional maps, we also test it with different sets of bases, such as Coordinates Manifold Harmonics (CMH) [73] and Orthogonalized eigenproducts (Orthoprods) [70]. In all cases, we align the bases using the method based on descriptors preservation [86] and ZoomOut [79]. For a fair comparison, we use the same bases for both the surface and the volumetric approaches.
@@
The results of our experiments are summarized in Tab. 2, where we show the dataset averaged geodesic error (AGE) for both frameworks and the percentage of single pairs for which our approach obtains a better AGE (Success Rate). We also report the slowdown factor induced by volumes and the ratio between the vertex count of the surface and volume meshes. Interestingly, our approach is consistently more accurate than its surface counterpart, suggesting that incorporating volumetric information into the matching process is always a rewarding choice when a volumetric discretization is available. The increased accuracy comes at the cost of additional computational, due to the introduction of the internal vertices. Inspecting the Slowdown columns in Tab. 2 we can see that there is a close relationship between the slowdown and the number of vertices introduced, suggesting that the volumetric pipeline is not significantly slower per se, but suffers from the overhead induced by the additional vertices needed to discretize the volume.
From the scale of the AGE, we can deduce that VOL is not a really challenging dataset for shape matching (watertight shapes, shared connectivity); thus, introducing volumetric information provides a real advantage only with more informative bases, such as Orthoprods. The accuracy of our approach seems related to the choice of the basis also when dealing with significantly more challenging datasets like SHREC’20. This is due to topological errors introduced by tetrahedralization. The topological changes reflect on the LBO eigenbasis and the Orthoprods basis, decreasing the accuracy. The CMH basis uses the vertex coordinates as a regularizer, compensating for the error. We provide a more detailed discussion in Appendix H.
6 Conclusions
@@
We presented the first volumetric functional maps pipeline, showcasing real applications that exploit discrete and continuous signal transfer for solid texturing, connectivity transfer, and segmentation transfer for medical applications, also improving the accuracy of surface matching tasks. Our practical contributions just scratch the surface of this problem and we foresee interesting avenues for future research.
@@
Our connectivity transfer shows great potential but is far from being perfect. Indeed, always achieving a fully injective map is technically possible, but considering a large enough portion of the spectrum may easily become computationally prohibitive. Devising smarter techniques to retain more precision for fixed spectrum size a research direction worth investigating.
The higher precision in shape matching tasks comes at the cost of an additional computational effort, because the most expensive steps in the pipeline (eigendecomposition and basis alignment) suffer from the increased vertex count. This limitation is intrinsic to our approach, but it could be mitigated by integrating scalable algorithms for computing and aligning bases specifically designed for volumes.
@@
A theoretical study of what properties are invariant under a volume functional map must be carried out. Despite our empirical studies show that volumetric maps are more accurate than their surface counterparts, a more profound theoretical analysis would provide useful insight for further developments.
Finally, we emphasize the importance of releasing new volumetric datasets to support ongoing research on volume mapping, and in particular on spectral volume methods. As discussed in Sec. 5.4, surface datasets that are well established in the functional map community hardly extend to volumes because of the difficulty to produce tetrahedral meshes of acceptable quality, due to topological and geometric noise, (e.g., large missing parts, overlaps, multiple connected components, non-manifold configurations and self-intersections). While robust volume meshing methods such as [43, 23] can handle intersections with ease, determining semantically plausible separations between the inside and outside portions of the volume in the presence of large gaps remains an open challenge. Recently, a large scale dataset to support ongoing research on piece-wise volume mappings to convex or star-shaped domains has been released [18]. Having similar contributions to support spectral methods would greatly help practitioners and ultimately foster more research in this emerging field.